# Programmable Cognitive Bias in Social Agents

### v1

@@ Abstract
Figure 1: Existing social simulation experiments often use implicit natural language descriptions to specify agent behaviors. However, we found that these specifications often resulted in inconsistent and unpredictable agent behaviors. For example, Ⓐ real-world economists are supposed to be less susceptible to the Framing Effect than the general population Thaler & Sunstein (2008); however, Ⓑ agents based on implicit natural language specifications often produce inconsistent behaviors across models, and the expected differences in behavior across roles (e.g., economists being less prone than laypeople) are not reliably observed. To tackle this challenge, Ⓒ we introduce CoBRA  which enables researchers to explicitly specify the cognitive biases of LLM-based agents quantitatively, thereby producing precise and consistent behaviors across models.
@@ 1 Introduction
Figure 2: Example closed-loop workflow of CoBRA. A social scientist aims to create an agent with a moderate framing effect (e.g., 2.6 on a 0–4 scale). ① She specifies the desired bias level in CoBRA alongside the natural language agent description. ② CoBRA measures the agent’s framing effect using validated classical social science experiments (e.g., the Asian Disease study Tversky & Kahneman (1981)). ③ If the measured bias deviates from the specification, the Behavioral Regulation Engine iteratively adjusts the agent—through prompt engineering, activation modifications, or fine-tuning—until the agent reliably demonstrates the target bias.
@@
The simulation of complex human behaviors through LLM-based agents is rapidly emerging as a powerful tool for social scientists Bail (2024); Park et al. (2023), offering new ways to test social theories Anthis et al. (2025a); Wu et al. (2025) and explore sensitive or hard-to-study behaviors without real-world consequences Hwang et al. (2025). A critical component of these simulations is specifying agent behaviors Yang et al. (2025); Chandramouli et al. (2024); Goyal et al. (2024) to control how the agent responds to prompts. For example, Park et al. specified the attributes of one of the agents in their paper using the following natural language description: ”John Lin is a pharmacy shopkeeper at … who loves to help people. He is always looking for ways to make the process of getting medication easier for his customers…” Park et al. (2023).
@@
These implicit natural language descriptions have become the dominant approach for specifying social agent behaviors and are used in nearly all social simulation experiments Rath et al. (2025); Park et al. (2023); Zhou et al. (2024); Li et al. (2023a); Zheng et al. (2023); Chen et al. (2024b); Jia et al. (2024). In the descriptions, researchers often specify demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five Gosling et al. (2003) or MBTI Furnham (1996)), personal interests (e.g., hobbies, preferences), social relationships (e.g., friends, family, colleagues), and professional or task-oriented identities (e.g., judge, programmer). These practices rest on a growing understanding that LLMs, having been trained on vast text corpora, possess both perceptual skills Wang et al. (2024b); Liu et al. (2024) that enable them to detect subtle cues in implicit language Gallegos et al. (2024), and strong role-playing capabilities Park et al. (2023)—drawing on patterns that link behaviors to human characteristics Fast et al. (2016); Shao et al. (2023b).
@@
However, the research community has little empirical understanding of whether this approach is sufficient for reliable social simulations. For example, would the conclusion of the social simulation change if we change the foundation models, the prompts, or the temperatures? As a consequence, the conclusions of most LLM-based social simulations often are broad qualitative impressions—such as noting that a simulation is plausible and exhibits an echo chamber similar to the real world Wang et al. (2024a); Zheng & Tang (2024).
@@
In this paper, we first conducted a pilot experiment to empirically examine the limitations of implicit natural language specification (Section 3). We designed four agent specifications by adapting examples from prior work Chen et al. (2024b); Park et al. (2023), implemented them across four foundation models, and evaluated their responses using a classic social science experiment, the Asian Disease study Tversky & Kahneman (1981). The results show that (1) the same specification produced inconsistent behaviors across models, and (2) implicit specification did not reliably produce the expected behaviors (e.g., trained experts like economists are less prone to the framing effect List (2003)).
@@
We then developed CoBRA—the Cognitive Bias Regulator for Social Agents (Figure 1)—a novel toolkit that enables users to precisely program the cognitive biases exhibited by social agents across different models. Compared to implicit natural language specifications, CoBRA offers two key advantages. First, it enhances reproducibility of agent behaviors across models, which is critical given that access to specific foundation models may change over time Chen et al. (2024a). Second, it provides explicit and quantitative control over agent behaviors, rather than relying on unpredictable model behaviors Sclar et al. (2023). For example, a scientist can directly specify that an agent should be less susceptible to framing effects.
@@
At the heart of CoBRA is a novel closed-loop system primitive that continuously measures and regulates agents’ cognitive biases by grounding their behaviors in classic social science experiments. Classic experiments provide validated and structured protocols: they define precise stimuli, expected behavioral distributions, and well-documented bias effects. For instance, the Asian Disease study Tversky & Kahneman (1981) has been extensively replicated and is widely regarded as a canonical demonstration of the framing effect: when identical outcomes are framed in terms of lives saved versus lives lost, human participants systematically shift between risk-averse and risk-seeking choices. Within CoBRA, such experiments function as calibration tasks: the system first measures whether an agent demonstrates the expected bias by observing its responses in these experiments and then adjusts the agent’s specification to align it with the target behavioral profile (Figure 2). We further hypothesize that aligning an agent’s behavior with target profiles in a set of classical experiments will reliably induce the corresponding cognitive biases, which will then generalize and manifest across a broader range of tasks (Section 7.2).
@@
Our implementation of CoBRA has two components: (1) Cognitive Bias Index that measures the cognitive bias of a social agent by quantifying the agent’s reactions in a set of validated classical social science experiments; (2) Behavioral Regulation Engine that aligns the agent’s behavior to demonstrate controlled cognitive bias. The Cognitive Bias Index contains eight classic social experiments mapped to four representative biases—Authority Effect, Bandwagon Effect, Confirmation Bias, and Framing Effect. The Behavioral Regulation Engine employs three approaches (prompt engineering Rath et al. (2025); Zhuo et al. (2024), representation engineering Zou et al. (2023); Rimsky et al. (2024); Todd et al. (2024), and fine-tuning Ouyang et al. (2022); Rafailov et al. (2023); Hu et al. (2022); Ilharco et al. (2023)) to align agent behaviors.
@@
We position CoBRA as an HCI toolkit Nebeling (2017) that empowers social scientists to create new types of social simulations. We evaluated CoBRA using two common strategies for assessing HCI toolkits Ledo et al. (2018): Technical Benchmark and Demonstration. In the benchmark, we show that (1) the cognitive bias induced through one classical experiment generalizes to other tasks (Section 7.2); (2) agents specified through CoBRA exhibit consistent behavioral tendencies regardless of the foundation model, sampling temperature, or reasoning mode (Section 7.3); and (3) CoBRA enables precise control under both open-weight and API-only settings (Section 7.4). In the Demonstration, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment Kramer et al. (2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated (Section 8). The simulated experiment validates a classical social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
@@ 2 Background & Related Work
This section reviews the use of large language models (LLMs) for social simulation and outlines the prevailing methodological pipeline, noting that current implicit natural-language agent specification approaches—while expressive and flexible—hinder rigor and comparability Bail (2024); Pang et al. (2025).
@@ 2.1 Believable Agents for Social Simulation
Believable Agents are interactive computational entities, originating as non-player characters in games Bates (1994). Early designs relied on rule-based systems or behavior trees to dictate their actions and decision-making processes Loyall (1997). Their observable behavior appears coherent, context-appropriate, and consistent with prior experiences, creating an “illusion of life” or human-like plausibility Park et al. (2023); Anthis et al. (2025b); Banovic et al. (2023); Ehsan et al. (2024). The goal is not human-equivalent cognition but externally credible behavior: agents should perceive, plan, and act in ways that are explainable, temporally consistent, and aligned with social and physical norms Goyal et al. (2024); Chandramouli et al. (2024); Liu et al. (2025a); Zhang et al. (2023).
Recent works use LLMs to create Believable Agents for social simulation Bail (2024); Wu et al. (2022); Anthis et al. (2025a); Hwang et al. (2025); Tjuatja et al. (2024); Prpa et al. (2024); Aubin Le Qu’er’e et al. (2024). For example, prior work Park et al. (2023); Zhou et al. (2024) utilizes LLMs’ language understanding to prompt models with natural language persona descriptions, creating LLM-based agents that interact within simulated environments. These agents facilitate the study of emergent phenomena—such as cooperation, norm formation, and conflict—and heuristic behaviors like conformity Liu et al. (2025b); Zhang et al. (2024); Shao et al. (2023a); Wu et al. (2025); Zhu et al. (2025). While these LLM-based Believable Agents broaden feasible questions and accelerate iteration by reducing labor, and mitigating ethical concerns Bail (2024); Hwang et al. (2025), they struggle with consistent, controllable behavior Yang et al. (2025); Han et al. (2025)—a critical need for rigorous social science simulations Bail (2024).
@@ 2.2 Reproducibility of LLM-Based Methods in Social Science
Reproducibility is a critical concern in the use of LLMs for social science research, as the probabilistic nature of LLMs combined with varying implementation methods often leads to inconsistent and non-reproducible outcomes Bail (2024); Zhao et al. (2021).
One major issue is that different LLMs exhibit divergent behaviors even with the same prompts Liu et al. (2023). Variations in training datasets Tayo et al. (2023), model architectures Jiang et al. (2024), and fine-tuning processes Touvron et al. (2023) exacerbate this inconsistency, while the opacity of proprietary models further complicates cross-model comparisons and limits generalizability Gudibande et al. (2023). Second, inference-time configurations significantly affect reproducibility. Sampling parameters like temperature settings influence output variability, with models differing in how peaky or flat their probability distributions are, affecting the generation coherence and diversity Peeperkorn et al. (2024). Similarly, reasoning modes Wei et al. (2022) (direct response vs. multi-step reasoning) introduce additional variability during inference, as different reasoning strategies can lead to divergent conclusions from the same input. Finally, post-training methods (e.g., RLHF Ouyang et al. (2022)) also introduce inconsistencies across models, further complicating reproducibility. Addressing these challenges requires standardized methodologies and explicit control mechanisms to ensure reliable results.
@@ 2.3 Classic Social Experiments
Classic social experiments French et al. (1959); Asch (1951); Tversky & Kahneman (1981); Milgram (1963); Kahneman & Tversky (2013); Wason (1960); Lord et al. (1979); Goldstein et al. (2008); Milgram et al. (1969) provide structured and validated knowledge: long-replicated classic social experimental paradigms operationalize social-psychological constructs into clear manipulations and measurement protocols Likert (1932); Vazire & Mehl (2008); Lilienfeld et al. (2017); the resulting behavioral patterns have been extensively validated and distilled into reusable, codified behavioral patterns Unknown (2015); Baron (2007); Cialdini (1993); Thaler & Sunstein (2008).
As a concrete focus in this paper, we ground agent specification in cognitive biases Unknown (2015); Baron (2007); Sumita et al. (2025)—a representative class of codified behavioral patterns—defined as systematic, predictable deviations from normative judgment and decision-making that arise from heuristics Tversky & Kahneman (1974); Simon (1990). Prominent examples include Confirmation Bias (favoring information that confirms pre-existing beliefs, validated through the Wason Selection task Wason (1960) and Biased Information study Lord et al. (1979)), Bandwagon Effect (adopting behaviors due to social conformity, demonstrated in Asch’s Line experiment Asch (1951) and the Hotel Towel study Goldstein et al. (2008)), and so on. These cognitive biases, repeatedly validated in classic experiments, provide a rigorous foundation for specifying and evaluating LLM-based agent behaviors.
@@ 3.1 Method
Our key idea is to use a classical social experiment, the Asian Disease study Tversky & Kahneman (1981), as the study instrument. We created agents using different specification methods, powered these agents using different foundation models, and then analyzed their responses in the simulated social experiment.
Agent specifications. We evaluated two representative examples of implicit natural language agent specifications: a persona-based specification from Park et al. (2023), which constructs agents with human-like profiles by incorporating demographics, personality traits, interests, and social relationships; and a role-based specification from Chen et al. (2024b), which defines agents through task-oriented identities, assigning roles, domain expertise, and interaction protocols to guide task-specific behavior.
For the persona-based example Park et al. (2023), we introduced three conditions: Common (an ordinary person), Economist (a persona with a professional background expected to make them less susceptible to framing effects and, therefore, less likely to prefer the positive frame compared to the Common persona), and Blank (no specification). We adopted the John Lin prompt from Park et al. (2023) to represent the Common persona (see Appendix B for full prompt):
@@
For the role-based example Chen et al. (2024b), we introduced two conditions: Teacher (the agent was specified as having the profession of a teacher) and Blank (no specification). We adopted the Teacher prompt from Chen et al. (2024b) with the Problem filled using the Asian Disease scenario prompt and the Student’s Solution presented in either a positive or negative frame (see Appendix B for full prompt):
@@
Simulated social experiment setup. We adapted the classic Asian Disease study Tversky & Kahneman (1981) for testing framing effect-participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 people will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 people will die”), into a prompt-based scenarios template. A sample scenario prompt is as follows:
@@
Figure 3: The example of persona-based specification Park et al. (2023) failed to produce consistent agent behavior regarding the framing effect in the Asian Disease paradigm. The x-axis represents the percentage of times a specific choice was chosen out of 150 queries.
Figure 4: The example of role-based specification Chen et al. (2024b) failed to produce consistent agent behavior regarding the framing effect in the Asian Disease paradigm. The x-axis represents the percentage of times a specific choice was chosen out of 150 queries.

### v2

@@ Abstract
Figure 1: Existing social simulation experiments often use implicit natural language descriptions to specify agent behaviors. However, we found that these specifications often resulted in inconsistent and unpredictable agent behaviors. For example, Ⓐ real-world economists are supposed to be less susceptible to the Framing Effect than the general population (Thaler & Sunstein, 2008); however, Ⓑ agents based on implicit natural language specifications often produce inconsistent behaviors across models, and the expected differences in behavior across roles (e.g., economists being less prone than laypeople) are not reliably observed. To tackle this challenge, Ⓒ we introduce CoBRA  which enables researchers to explicitly specify the cognitive biases of LLM-based agents quantitatively, thereby producing precise and consistent behaviors across models.
@@ 1 Introduction
Figure 2: Example closed-loop workflow of CoBRA. A social scientist aims to create an agent with a moderate framing effect (e.g., 2.6 on a 0–4 scale). ① She specifies the desired bias level in CoBRA alongside the natural language agent description. ② CoBRA measures the agent’s framing effect using validated classical social science experiments (e.g., the Asian Disease study (Tversky & Kahneman, 1981)). ③ If the measured bias deviates from the specification, the Behavioral Regulation Engine iteratively adjusts the agent—through prompt engineering, activation modifications, or fine-tuning—until the agent reliably demonstrates the target bias.
@@
The simulation of complex human behaviors through LLM-based agents is rapidly emerging as a powerful tool for social scientists (Bail, 2024; Park et al., 2023), offering new ways to test social theories (Anthis et al., 2025a; Wu et al., 2025) and explore sensitive or hard-to-study behaviors without real-world consequences (Hwang et al., 2025). A critical component of these simulations is specifying agent behaviors (Yang et al., 2025; Chandramouli et al., 2024; Goyal et al., 2024) to control how the agent responds to prompts. For example, Park et al. specified the attributes of one of the agents in their paper using the following natural language description: ”John Lin is a pharmacy shopkeeper at … who loves to help people. He is always looking for ways to make the process of getting medication easier for his customers…” (Park et al., 2023).
@@
These implicit natural language descriptions have become the dominant approach for specifying social agent behaviors and are used in nearly all social simulation experiments (Rath et al., 2025; Park et al., 2023; Zhou et al., 2024; Li et al., 2023a; Zheng et al., 2023; Chen et al., 2024b; Jia et al., 2024). In the descriptions, researchers often specify demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five (Gosling et al., 2003) or MBTI (Furnham, 1996)), personal interests (e.g., hobbies, preferences), social relationships (e.g., friends, family, colleagues), and professional or task-oriented identities (e.g., judge, programmer). These practices rest on a growing understanding that LLMs, having been trained on vast text corpora, possess both perceptual skills (Wang et al., 2024b; Liu et al., 2024) that enable them to detect subtle cues in implicit language (Gallegos et al., 2024), and strong role-playing capabilities (Park et al., 2023)—drawing on patterns that link behaviors to human characteristics (Fast et al., 2016; Shao et al., 2023b).
@@
However, the research community has little empirical understanding of whether this approach is sufficient for reliable social simulations. For example, would the conclusion of the social simulation change if we change the foundation models, the prompts, or the temperatures? As a consequence, the conclusions of most LLM-based social simulations often are broad qualitative impressions—such as noting that a simulation is plausible and exhibits an echo chamber similar to the real world (Wang et al., 2024a; Zheng & Tang, 2024).
@@
In this paper, we first conducted a pilot experiment to empirically examine the limitations of implicit natural language specification (Section 3). We designed four agent specifications by adapting examples from prior work (Chen et al., 2024b; Park et al., 2023), implemented them across four foundation models, and evaluated their responses using a classic social science experiment, the Asian Disease study (Tversky & Kahneman, 1981). The results show that (1) the same specification produced inconsistent behaviors across models, and (2) implicit specification did not reliably produce the expected behaviors (e.g., trained experts like economists are less prone to the framing effect (List, 2003)).
@@
We then developed CoBRA—the Cognitive Bias Regulator for Social Agents (Figure 1)—a novel toolkit that enables users to precisely program the cognitive biases exhibited by social agents across different models. Compared to implicit natural language specifications, CoBRA offers two key advantages. First, it enhances reproducibility of agent behaviors across models, which is critical given that access to specific foundation models may change over time (Chen et al., 2024a). Second, it provides explicit and quantitative control over agent behaviors, rather than relying on unpredictable model behaviors (Sclar et al., 2023). For example, a scientist can directly specify that an agent should be less susceptible to framing effects.
@@
At the heart of CoBRA is a novel closed-loop system primitive that continuously measures and regulates agents’ cognitive biases by grounding their behaviors in classic social science experiments. Classic experiments provide validated and structured protocols: they define precise stimuli, expected behavioral distributions, and well-documented bias effects. For instance, the Asian Disease study (Tversky & Kahneman, 1981) has been extensively replicated and is widely regarded as a canonical demonstration of the framing effect: when identical outcomes are framed in terms of lives saved versus lives lost, human participants systematically shift between risk-averse and risk-seeking choices. Within CoBRA, such experiments function as calibration tasks: the system first measures whether an agent demonstrates the expected bias by observing its responses in these experiments and then adjusts the agent’s specification to align it with the target behavioral profile (Figure 2). We further hypothesize that aligning an agent’s behavior with target profiles in a set of classical experiments will reliably induce the corresponding cognitive biases, which will then generalize and manifest across a broader range of tasks (Section 7.2).
@@
Our implementation of CoBRA has two components: (1) Cognitive Bias Index that measures the cognitive bias of a social agent by quantifying the agent’s reactions in a set of validated classical social science experiments; (2) Behavioral Regulation Engine that aligns the agent’s behavior to demonstrate controlled cognitive bias. The Cognitive Bias Index contains eight classic social experiments mapped to four representative biases—Authority Effect, Bandwagon Effect, Confirmation Bias, and Framing Effect. The Behavioral Regulation Engine employs three approaches (prompt engineering (Rath et al., 2025; Zhuo et al., 2024), representation engineering (Zou et al., 2023; Rimsky et al., 2024; Todd et al., 2024), and fine-tuning (Ouyang et al., 2022; Rafailov et al., 2023; Hu et al., 2022; Ilharco et al., 2023)) to align agent behaviors.
@@
We position CoBRA as an HCI toolkit (Nebeling, 2017) that empowers social scientists to create new types of social simulations. We evaluated CoBRA using two common strategies for assessing HCI toolkits (Ledo et al., 2018): Technical Benchmark and Demonstration. In the benchmark, we show that (1) the cognitive bias induced through one classical experiment generalizes to other tasks (Section 7.2); (2) agents specified through CoBRA exhibit consistent behavioral tendencies regardless of the foundation model, sampling temperature, or reasoning mode (Section 7.3); and (3) CoBRA enables precise control under both open-weight and API-only settings (Section 7.4). In the Demonstration, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment (Kramer et al., 2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated (Section 8). The simulated experiment validates a classical social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
@@ 2 Background & Related Work
This section reviews the use of large language models (LLMs) for social simulation and outlines the prevailing methodological pipeline, noting that current implicit natural-language agent specification approaches—while expressive and flexible—hinder rigor and comparability (Bail, 2024; Pang et al., 2025).
@@ 2.1 Believable Agents for Social Simulation
Believable Agents are interactive computational entities, originating as non-player characters in games (Bates, 1994). Early designs relied on rule-based systems or behavior trees to dictate their actions and decision-making processes (Loyall, 1997). Their observable behavior appears coherent, context-appropriate, and consistent with prior experiences, creating an “illusion of life” or human-like plausibility (Park et al., 2023; Anthis et al., 2025b; Banovic et al., 2023; Ehsan et al., 2024). The goal is not human-equivalent cognition but externally credible behavior: agents should perceive, plan, and act in ways that are explainable, temporally consistent, and aligned with social and physical norms (Goyal et al., 2024; Chandramouli et al., 2024; Liu et al., 2025a; Zhang et al., 2023).
Recent works use LLMs to create Believable Agents for social simulation (Bail, 2024; Wu et al., 2022; Anthis et al., 2025a; Hwang et al., 2025; Tjuatja et al., 2024; Prpa et al., 2024; Aubin Le Qu’er’e et al., 2024). For example, prior work (Park et al., 2023; Zhou et al., 2024) utilizes LLMs’ language understanding to prompt models with natural language persona descriptions, creating LLM-based agents that interact within simulated environments. These agents facilitate the study of emergent phenomena—such as cooperation, norm formation, and conflict—and heuristic behaviors like conformity (Liu et al., 2025b; Zhang et al., 2024; Shao et al., 2023a; Wu et al., 2025; Zhu et al., 2025). While these LLM-based Believable Agents broaden feasible questions and accelerate iteration by reducing labor, and mitigating ethical concerns (Bail, 2024; Hwang et al., 2025), they struggle with consistent, controllable behavior (Yang et al., 2025; Han et al., 2025)—a critical need for rigorous social science simulations (Bail, 2024).
@@ 2.2 Reproducibility of LLM-Based Methods in Social Science
Reproducibility is a critical concern in the use of LLMs for social science research, as the probabilistic nature of LLMs combined with varying implementation methods often leads to inconsistent and non-reproducible outcomes (Bail, 2024; Zhao et al., 2021).
One major issue is that different LLMs exhibit divergent behaviors even with the same prompts (Liu et al., 2023). Variations in training datasets (Tayo et al., 2023), model architectures (Jiang et al., 2024), and fine-tuning processes (Touvron et al., 2023) exacerbate this inconsistency, while the opacity of proprietary models further complicates cross-model comparisons and limits generalizability (Gudibande et al., 2023). Second, inference-time configurations significantly affect reproducibility. Sampling parameters like temperature settings influence output variability, with models differing in how peaky or flat their probability distributions are, affecting the generation coherence and diversity (Peeperkorn et al., 2024). Similarly, reasoning modes (Wei et al., 2022) (direct response vs. multi-step reasoning) introduce additional variability during inference, as different reasoning strategies can lead to divergent conclusions from the same input. Finally, post-training methods (e.g., RLHF (Ouyang et al., 2022)) also introduce inconsistencies across models, further complicating reproducibility. Addressing these challenges requires standardized methodologies and explicit control mechanisms to ensure reliable results.
@@ 2.3 Classic Social Experiments
Classic social experiments (French et al., 1959; Asch, 1951; Tversky & Kahneman, 1981; Milgram, 1963; Kahneman & Tversky, 2013; Wason, 1960; Lord et al., 1979; Goldstein et al., 2008; Milgram et al., 1969) provide structured and validated knowledge: long-replicated classic social experimental paradigms operationalize social-psychological constructs into clear manipulations and measurement protocols (Likert, 1932; Vazire & Mehl, 2008; Lilienfeld et al., 2017); the resulting behavioral patterns have been extensively validated and distilled into reusable, codified behavioral patterns (Unknown, 2015; Baron, 2007; Cialdini, 1993; Thaler & Sunstein, 2008).
As a concrete focus in this paper, we ground agent specification in cognitive biases (Unknown, 2015; Baron, 2007; Sumita et al., 2025)—a representative class of codified behavioral patterns—defined as systematic, predictable deviations from normative judgment and decision-making that arise from heuristics (Tversky & Kahneman, 1974; Simon, 1990). Prominent examples include Confirmation Bias (favoring information that confirms pre-existing beliefs, validated through the Wason Selection task (Wason, 1960) and Biased Information study (Lord et al., 1979)), Bandwagon Effect (adopting behaviors due to social conformity, demonstrated in Asch’s Line experiment (Asch, 1951) and the Hotel Towel study (Goldstein et al., 2008)), and so on. These cognitive biases, repeatedly validated in classic experiments, provide a rigorous foundation for specifying and evaluating LLM-based agent behaviors.
@@ 3.1 Method
Our key idea is to use a classical social experiment, the Asian Disease study (Tversky & Kahneman, 1981), as the study instrument. We created agents using different specification methods, powered these agents using different foundation models, and then analyzed their responses in the simulated social experiment.
Agent specifications. We evaluated two representative examples of implicit natural language agent specifications: a persona-based specification from (Park et al., 2023), which constructs agents with human-like profiles by incorporating demographics, personality traits, interests, and social relationships; and a role-based specification from (Chen et al., 2024b), which defines agents through task-oriented identities, assigning roles, domain expertise, and interaction protocols to guide task-specific behavior.
For the persona-based example (Park et al., 2023), we introduced three conditions: Common (an ordinary person), Economist (a persona with a professional background expected to make them less susceptible to framing effects and, therefore, less likely to prefer the positive frame compared to the Common persona), and Blank (no specification). We adopted the John Lin prompt from (Park et al., 2023) to represent the Common persona (see Appendix B for full prompt):
@@
For the role-based example (Chen et al., 2024b), we introduced two conditions: Teacher (the agent was specified as having the profession of a teacher) and Blank (no specification). We adopted the Teacher prompt from (Chen et al., 2024b) with the Problem filled using the Asian Disease scenario prompt and the Student’s Solution presented in either a positive or negative frame (see Appendix B for full prompt):
@@
Simulated social experiment setup. We adapted the classic Asian Disease study (Tversky & Kahneman, 1981) for testing framing effect-participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 people will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 people will die”), into a prompt-based scenarios template. A sample scenario prompt is as follows:
@@
Figure 3: The example of persona-based specification (Park et al., 2023) failed to produce consistent agent behavior regarding the framing effect in the Asian Disease paradigm. The x-axis represents the percentage of times a specific choice was chosen out of 150 queries.
Figure 4: The example of role-based specification (Chen et al., 2024b) failed to produce consistent agent behavior regarding the framing effect in the Asian Disease paradigm. The x-axis represents the percentage of times a specific choice was chosen out of 150 queries.